[Hallc_running] [New Logentry] Hall C update - ESR trip

gaskelld at jlab.org gaskelld at jlab.org
Wed Dec 6 17:55:02 EST 2017

Logentry Text:
The ESR tripped around 3 pm today. The Hall C hydrogen target warmed up, and the SHMS and HMS magnet helium reservoirs are empty.  During ESR recovery, the magnets in the halls have priority - the soonest the Hall C target would be cooled down is tomorrow (Thursday).

Tonight: Wednesday swing, and Thursday owl target shifts are not required.  The Thursday, Day shift target operator should come in as scheduled. Shift leader and 3rd's are still on call until day shift, Friday.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3500537
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171206/284ec35f/attachment.html>

More information about the Hallc_running mailing list